OpenStudy (anonymous):

Find the solution of the given initial value problem in explicit form. y' = (1-11x)y^2, y(0) = -1/8

OpenStudy (anonymous):

this is saperable

zepdrix (zepdrix):

\[\Large\bf\sf y'\quad=\quad (1-11x)y^2\]Hmm so this appears to be separable.\[\Large\bf\sf \frac{dy}{y^2}\quad=\quad (1-11x)dx\]From here we integrate. Understand that much of it?

OpenStudy (anonymous):

oh ok i got -y^-1 = x-(11/2)^2 is that right?

zepdrix (zepdrix):

(11/2)x^2? Mmm ya looks good! Don't forget a constant of integration somewhere\[\Large\bf\sf -\frac{1}{y}\quad=\quad x-\frac{11}{2}x^2+c\]

zepdrix (zepdrix):

So we need to solve for y as our next step.

zepdrix (zepdrix):

err actually, let's use the initial data before solving for y explicitly. It'll be easier from here I think.

zepdrix (zepdrix):

\[\Large\bf\sf y(0)=-\frac{1}{8}\]Plugging this in:\[\Large\bf\sf 8\quad=\quad 0-\frac{11}{2}0^2+c\]Understand how that left side simplified?

OpenStudy (anonymous):

hmmm... how did -1/y become 8?

OpenStudy (anonymous):

ohhh never mind i got it.